OpenStudy (anonymous):

reduce to lowest terms (38x^2yz^2)/(-19xy^2z^3)

OpenStudy (anonymous):

38/-19=-2 \[\frac{x^2}{x}=x\] \[\frac{y}{y^2}=\frac{1}{y}\] \[\frac{z^2}{z^3}=\frac{1}{z}\] multiplying together, we get \[\frac{-2x}{yz}\]